# Frank–Wolfe algorithm for DC optimization problem

### v2

@@ 7 Conclusions
We analyzed Algorithm 1 and its finite–difference variant Algorithm 2 within the DC piecewise-star-convex framework. For piecewise-star-convex objectives, the exact–gradient method attains the convex–case rate $\mathcal{O}(1/k)$ for both the function suboptimality $f(x_{k})-f^{\star}$ and the optimality certificate (duality/Frank–Wolfe gap). For the finite–difference method, we established asymptotic convergence and an $\mathcal{O}(1/\sqrt{k})$ decay of the gap. A full iteration–complexity theory for the finite–difference variant under piecewise-star-convex remains open. Promising directions include: (i) matching $\mathcal{O}(1/k)$ rates via tighter control of finite–difference bias and adaptive radii, (ii) guarantees that account for possible transitions across piecewise-star-convex cells without stabilization assumptions, and (iii) extensions to stochastic oracles, away–steps, and line–search variants.

### v3

@@ 6 Conclusions
We analyzed Algorithm 1 within the DC piecewise–star–convex framework. In particular, for functions that are piecewise–star–convex on a compact convex set, we showed that the proposed Frank–Wolfe scheme equipped with the backtracking stepsize attains the same sublinear iteration complexity as in the convex case, namely, the function-value error (measured against the appropriate cellwise minimum) and the Frank–Wolfe gap both decay on the order of $1/k$. These results extend convex-type complexity guarantees to a broad class of nonconvex and potentially nonsmooth DC objectives, complementing recent rates established for globally star–convex functions. Natural directions for further work include, identifying verifiable conditions under which the iterates enter (and remain in) a single cell so that the bounds hold from the first iteration, refining the analysis to better quantify the transient phase before the “active” cell is identified, and extending the guarantees to common Frank–Wolfe enhancements such as away–steps and pairwise steps, as well as other line-search mechanisms.